OpenStudy (bibby):

(c) Since π/4 = arctan, you can find the value of π/4 using your expansion. How many terms of the expansion would you have to add in order to obtain π correct to four decimal places? edit: forgot d and e... (d) In an earlier assignment you found the Euler-Machin identity \[\pi/4 = 4\arctan \frac{ 1 }{ 5 } - \arctan \frac{ 1 }{ 239 }\] Using your series for arctan x to evaluate arctan 1/5 and arctan 1/239, how many terms would you have to find pi correct to 4 decimal places? (e) find pi to 4 decimal places using the Maclaurin series for arctan x done parts a and b (a) Find the Maclaurin series for arctan(x) (b) Show that the Maclaurin series for arctan(x) converges on x=1

OpenStudy (shamil98):

for a \[\huge \sum_{n=0}^{\infty} \frac{ (-1)^nx^{2n+1} }{ 2n+1 }\]

OpenStudy (shamil98):

google is great :D

OpenStudy (bibby):

lol, I did part a by successive derivation

OpenStudy (bibby):

Not sure how to approach b. For part c, since it's alternating, does approximation go by the next term you don't count in the series?

OpenStudy (shamil98):

for b, i think you just need to use the summation series , and input x = 1

OpenStudy (shamil98):

im prob wrong, i'd ask the other wizards of math on here >.<

OpenStudy (bibby):

Wizards of OS lend me your energy :* @Hero @ranga @nincompoop @shamil98 @bibby

OpenStudy (shamil98):

d/dx arctan = 1/(1+x^2) meh ik dat much o.o

OpenStudy (the_fizicx99):

Ask @math&ing001 ?? Idk :<

OpenStudy (bibby):

\[\pi/4 = 4\arctan \frac{ 1 }{ 5 } - \arctan \frac{ 1 }{ 239 }\] I meant

OpenStudy (anonymous):

For (a), there's a rather nice way of getting the Maclaurin series for \(\large \arctan x\). Recall that \[\large \frac{d}{dx}\arctan x= \frac{1}{1+x^2}\] So, if we can come up with a Maclaurin series for \(\large \dfrac{1}{1+x^2}\) and then integrate termwise, we've got our series for \(\large \arctan x\). Now, what's the series representation for \(\large\dfrac{1}{1+x^2}\)? Well, we know for \(\large |x|<1\) that \[1+x+x^2+x^3+\ldots = \sum_{n=0}^{\infty}x^n=\large\frac{1}{1-x}\](this is a geometric series, by the way). So, we now see that if \(\large |x|<1\), then clearly \(\large |-x^2|=|x^2|<1\) and thus \[\large\frac{1}{1+x^2} = \frac{1}{1-(-x^2)} = \sum_{n=0}^{\infty} (-x^2)^n = \sum_{n=0}^{\infty}(-1)^nx^{2n}\] Therefore, we see that \[\large \begin{aligned}\arctan x &= \int\frac{1}{1+x^2}\,dx \\ &= \int\left(\sum_{n=0}^{\infty}(-1)^nx^{2n}\right)\,dx \\ &= C + \sum_{n=0}^{\infty}\left((-1)^ n\int x^{2n}\,dx\right)\\ &= C+\sum_{n=0}^{\infty} \frac{x^{2n+1}}{2n+1}\end{aligned}\] Now, we know that \(\large \arctan 0=0\); thus \[\large 0=\arctan 0 = C+\sum_{n=0}^{\infty} \frac{0^{2n+1}}{2n+1}=C \implies C=0\] Therefore, the Maclaurin series for arctan x is \[\large \arctan x=\sum_{n=0}^{\infty} \frac{x^{2n+1}}{2n+1}.\] ----------- For (b), when x=1, the series we found in (a) turns into \[\sum_{n=0}^{\infty}\frac{(-1)^n}{2n+1}\]You can then show that this series converges using the alternating series test. ---------- I will have to think about the other parts a little more before answering those, but I hope this other way of computing the Maclaurin series makes sense! :-)

OpenStudy (bibby):

Gosh. I'd write a verbose paragraph but all I can really say is thank you for putting in this much work. You're an amazing person. Now I must read and digest this :p

OpenStudy (anonymous):

Woops; the \((-1)^n\) term went missing in my post for some reason... sigh.

OpenStudy (anonymous):

If only there was a nice way to edit your post... >_>. The \(\large(−1)^n\) term should appear in the following lines: Where I say \(\large \displaystyle =C+\sum_{n=0}^{\infty}\frac{\color{red}{(−1)^ n}x^{2n+1}}{2n+1}\) in the integration steps. Where I say \(\large\displaystyle 0=\arctan 0=C+\sum_{n=0}^{\infty}\frac{\color{red}{(−1)^n}0^{2n+1}}{2n+1}=C\implies C=0\) And then where I officially claimed \(\large \displaystyle \arctan x=\sum_{n=0}^{\infty}\frac{\color{red}{(−1)^n} x^{2n+1}}{2n+1}\). Sorry about that! :-/

OpenStudy (anonymous):

lol

OpenStudy (bibby):

FYI: I have parts a and b.

OpenStudy (the_fizicx99):

I can explain this to you bibby :>

OpenStudy (bibby):

lmao I'll repost it. leave me alone

OpenStudy (the_fizicx99):

NO. I GOT THIS. SIT.

OpenStudy (bibby):

I'll be back in a bit bro :*

OpenStudy (the_fizicx99):

kay :[